TRUE/FALSE. Write T if the statement is true and F if the statement is false. 11) As long as income is above planned aggr

help
Answer #1

11) False. If income is greater than PAE, it means there is an unplanned inventory investment not reduction because demand is less than supply of goods and services

12) This is true.

13) False. Firms will decrease production if inventory is piling up

14) False, multiplier is 1/(1-MPC) = 1/(1-0.8) = 5

15) True

16) True

17) True

18) True

19) False. They will accumulate

20) false. It includes change in inventories